OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <8, 7>, v = <9, 7>

OpenStudy (anonymous):

Got it!

zepdrix (zepdrix):

Oh good job! c:

OpenStudy (anonymous):

Actually I need help :( @zepdrix

OpenStudy (anonymous):

I'm stuck in a part where I have (121)/ the square root of 145 and the square root of 98

zepdrix (zepdrix):

So how are we finding the angle between them? Using this identity?\[\large \cos \theta=\frac{\vec u \cdot \vec v}{|u|\cdot|v|}\]Hopefully I remembered that correctly...

OpenStudy (anonymous):